OpenStudy (anonymous):

Match each item in Column A to an answer in Column B. 1. (1 pt) Which functions have a positive rate of change and which have a negative rate of change? Column A Column B 1. A.positive 2x – 3y = 9 B.negative 2. y = −4x + 5 3. x y −1 −5 1 1 3 7 4. x y 1 −2 2 −6 3 −10

OpenStudy (anonymous):

@roz_mar23 @sleepyjess

OpenStudy (texaschic101):

rate of change is also the slope for #'s 1 and 2....put equation in y = mx + b form where m represents your slope. If the slope is positive, it has a positive rate...if negative, it has a negative rate. for #'s 3 and 4.....you will need to use slope formula slope = (y2 - y1) / (x2 - x1)......sub in 2 sets of the points given....and find the slope any questions ?

OpenStudy (anonymous):

@texaschic101 would you mind guideing me though this

OpenStudy (anonymous):

@texaschic101 ?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

@tylermcmullen23 @NeonStrawsForever

OpenStudy (texaschic101):

1. 2x - 3y = 9 (put in y = mx + b form) -3y = -2x + 9 y = 2/3x - 3......slope(m) = 2/3.......positive slope 2. y = -4x + 5 slope(m) = -4.......negative slope 3. (y2 - y1) / (x2 - x1) (-1,-5)....x1 = -1 and y1 = -5 (1,1)...x2 = 1 and y2 = 1 now sub (1 - (-5) / (1 - (-1) (1 + 5) / (1 + 1) 6/2 = 3.......so the slope is 3......positive slope 4. (y2 - y1) / (x2 - x1) (1,-2)...x1 = 1 and y1 = -2 (2,-6)...x2 = 2 and y2 = -6 now we sub (-6 - (-2) / (2 - 1) (-6 + 2) / 1 -4/1.....slope is -4.....negative slope any questions ?